OpenStudy (anonymous):

At the market you purchase some apples and oranges. There were 12 items purchased. You buy twice as many apples as oranges. a. Define the variables. b. Write an algebraic equation using one variable for the above situation. c. Solve this equation for the number of apples and oranges you purchase

OpenStudy (anonymous):

variables are what u want them to be a believe

OpenStudy (anonymous):

so x stands for oranges and y stands for apples

OpenStudy (texaschic101):

A. ) a = apples and o = oranges (be careful..that is o not a zero) B) a = 2o 2o + o = 12 <-- your equation with 1 variable C) 2o + o = 12 3o = 12 o = 12/3 o = 4 a + o = 12 a + 4 = 12 a = 12 - 4 a = 8 check.. 2o + o = 12 2(4) + 4 = 12 8 + 4 = 12 12 = 12 (correct) there are L 4 oranges and 8 apples

OpenStudy (anonymous):

u would get y+X+X =12

OpenStudy (texaschic101):

oops..there are 4 oranges and 8 apples

OpenStudy (anonymous):

yes indeed

OpenStudy (anonymous):

gg

OpenStudy (texaschic101):

it might be easier to use x and y instead of a and o....less confusing

OpenStudy (anonymous):

ok i have to use the variables given im in a timed exam

OpenStudy (anonymous):

might be

OpenStudy (anonymous):

the variables are what you want them to be most of the time so it doesnt really matter unless ^^ given to you

OpenStudy (texaschic101):

there are no variables given.....you can let a = apples..or you can let any letter = apples and any letter = oranges....it does not matter what letter you use

OpenStudy (texaschic101):

just be consistent with you variables (letters)

OpenStudy (anonymous):

oh lol i didnt even look

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

no not really im just gunna not be lazy and do it myself i was just looking for something to copy and paste

OpenStudy (texaschic101):

ok...well, if you need any more help, just tag me

OpenStudy (anonymous):

ok thanks